Question 4 Given vectors u = i + 3j and v = 21-5j, find 1.- v 2. u-v uz saved at
Answer #1

• given a = î+39, vzdî-55 0 find 3 J = 73 (21-59) = -3824 + 3x59 -352 +3 h o u_v= (û +32) - (2 1 5 h) - ti )+(345) qua 7 8